Solve the system of equations by substitution:
3x - y = 4
x + 5y = -4

Answer by lenny460(1073)

Solve the system of equations by substitution:
3x - y = 4 (Equation 1)
x + 5y = -4 (Equation 2)

3x - y = 4
Subtract 3x from each side of the equation.
3x - y - 3x = 4 - 3x
-y = 4 - 3x
Divide each side by -1
y = -4 + 3x (Equation 3)
Substitute Equation 3 into Equation 2.
x + 5y = -4
x + 5(-4 + 3x) = -4
x - 20 + 15x = -4
16x - 20 = -4
16x - 20 + 20 = -4 + 20
16x = 16
x = 1

Substitute x = 1 into Equation 1.

3x - y = 4
3(1) - y = 4
3 - y = 4
3 - y - 3 = 4 - 3
-y = 1
Divide each side by -1
y = -1


Answer:
x = 1
y = -1






lennox obuong
algebra tutor